# Fast Policy Learning for Linear-Quadratic Control with Entropy Regularization

### v3

@@ 3 Analysis of Value Function and Policy Gradient
A function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is considered smooth if the following condition is satisfied [12, 18]:
$\left|f(x)-f(y)+\nabla f(x)^{\top}(y-x)\right|\leq\frac{m}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{n},$
with $m$ a finite constant. In general, characterizing the smoothness of $C(K,\Sigma)$ is challenging, as it may become unbounded when the eigenvalues of $A-BK$ exceed $\frac{1}{\sqrt{\gamma}}$ or when $\sigma_{\min}(\Sigma)$ is close to 00. Nevertheless, in Lemma 3.3, we will see that if $C(K,\Sigma)$ is “almost” smooth, then the difference $C(K,\Sigma)-C(K^{\prime},\Sigma^{\prime})$ can be bounded by the sum of linear and quadratic terms involving $K-K^{\prime}$ and $\Sigma-\Sigma^{\prime}$.
@@ 4 Regularized Policy Gradient Method
RPG takes into account the inherent structure of the parameter space, which can accelerate convergence.
By the explicit expressions of $\nabla_{K}C(K,\Sigma)$ and $\nabla_{\Sigma}C(K,\Sigma)$ in (18), the above update is equivalent to
$K^{(t+1)}$ $=K^{(t)}-2\eta_{1}E_{K^{(t)}},$ (RPG)
@@
Assume that the covariance matrix $\Sigma$ is parameterized as scalar multiplication of an identity matrix, i.e., $\Sigma=sI$ for some $s>0$ and
$\pi_{K,s}(x,u)=\mathcal{N}(Kx,sI).$ (20)
Then the natural policy gradient follows the update [26]:
$K^{\prime}$ $=K-\eta G_{K}^{-1}\nabla C(K,sI),\,\,s^{\prime}=s-\eta G_{s}^{-1}\partial_{s}C(K,sI),$
where
$G_{K}=\mathbb{E}\left[\sum_{t=0}^{\infty}\,\nabla\log\pi_{K,s}(u_{t}|x_{t})\nabla\log\pi_{K,s}(u_{t}|x_{t})^{\top}\right],$ (21)
$G_{s}=\mathbb{E}\left[\sum_{t=0}^{\infty}\,\partial_{s}\log\pi_{K,s}(u_{t}|x_{t})\partial_{s}\log\pi_{K,s}(u_{t}|x_{t})^{\top}\right]$ (22)
is the Fischer information matrix under policy $\pi_{K,s}$. When the covariance matrix of the Gaussian policy takes a diagonal form as in (20), (4.1) and (22) are equivalent to
$K^{\prime}=K-\eta\nabla C(K,sI)(S_{K,sI})^{-1},\quad s^{\prime}=s-\tilde{\eta}~{}\partial_{s}C(K,sI)s^{2},$ (23)
@@
Even though (RPG) is similar to (23), there is no corresponding Fisher information form associated with (RPG) because of the additional step that simultaneously updates the covariance matrix $\Sigma$, which may not necessarily be diagonal.
@@
Given $\tau\in(0,2\sigma_{\min}(R)],$ take $(K^{(0)},\Sigma^{(0)})\in\Omega$ such that $\Sigma^{(0)}\preceq I$. Define $M_{\tau}:=\frac{\tau k}{2(1-\gamma)}\log\left(\frac{\sigma_{\min}(R)}{\pi\tau}\right)$ and
$r_{0}:=\max\left\{\frac{2}{\tau\sigma_{\min}\left(\Sigma^{(0)}\right)},\,\|R\|+\gamma\frac{\|B^{\top}B\|\left(C(K^{(0)},\Sigma^{(0)})-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}\right\}.$
Then for $\eta_{1},\eta_{2}=\frac{1}{2r_{0}},\frac{\tau(1-\gamma)}{2r_{0}^{2}}$ and for
$N\geq\max\left\{\frac{\|S_{K^{*},\Sigma^{*}}\|r_{0}}{\mu\sigma_{\min}(R)},\frac{8r_{0}^{3}}{\tau^{2}\sigma_{\min}(R)}\right\}\log\frac{C(K^{(0)},\Sigma^{(0)})-C(K^{*},\Sigma^{*})}{\varepsilon},$
the regularized policy gradient descent (RPG) has the following performance bound:
@@
To prove Theorem 4.2, we will first need the boundedness of the one-step update of the covariance $\Sigma$, in order to guarantee the well-definedness of the cost function along the trajectory when performing (RPG) (Lemma 4.3). Additionally, we will bound the one-step update of (RPG) (Lemma 4.4), and provide an upper bound of $\|P_{K}\|$ in terms of the objective function $C(K,\Sigma)$ (Lemma 4.5).
@@
[…]
@@
with $10<\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}<$ 1 .
@@
Let $M_{\tau}$ be defined in the same way as in Theorem 4.2. Then for all $(K,\Sigma)\in\Omega$,
$C(K,\Sigma)\geq\left(\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)\right)\|P_{K}\|+M_{\tau}.$
Proof.
@@
$\frac{1}{\left\|R+\gamma B^{\top}P_{K^{(t)}}B\right\|}\geq\frac{1}{\|R\|+\gamma\|B^{\top}B\|\left\|P_{K^{(t)}}\right\|}\geq\frac{1}{\|R\|+\gamma\frac{\|B^{\top}B\|\left(C({K^{(t)}},\Sigma^{(t)})-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}}.$ (24)
Let $a=\frac{\tau}{2r_{0}}$, $\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}$. The proof is completed by induction to show $C(K^{(t+1)},\Sigma^{(t+1)})\leq(1-\zeta)C(K^{(t)},\Sigma^{(t)})$ and $aI\preceq\Sigma^{(t+1)}\preceq I$ holds for all $t\geq 0$: at $t=0$, apply (24) to get $\eta_{1}\leq\frac{1}{2\|R+\gamma B^{\top}P_{K^{(0)}}B\|}$ and $a\leq\frac{\tau}{2\|R+\gamma B^{\top}P_{K^{(0)}}B\|}$. Additionally with $\eta_{2}=\frac{\tau(1-\gamma)}{2r_{0}^{2}}=\frac{2(1-\gamma)a^{2}}{\tau}$ and $aI\preceq\Sigma^{(0)}\preceq I$, Lemma 4.4 can be applied to get
$C(K^{(1)},\Sigma^{(1)})\leq(1-\zeta)C(K^{(0)},\Sigma^{(0)})\leq C(K^{(0)},\Sigma^{(0)}),$
and $aI\preceq\Sigma^{(1)}\preceq I$. The proof proceeds by arguing that Lemma 4.4 can be applied at every step. If it were the case that $C\left(K^{(t)},\Sigma^{(t)}\right)\leq(1-\zeta)C\left(K^{(t-1)},\Sigma^{(t-1)}\right)\leq C\left(K^{(0)},\Sigma^{(0)}\right)$ and $aI\preceq\Sigma^{(t)}\preceq I$, then
$2\eta_{1}$ $=\frac{1}{\|R\|+\gamma\frac{\|B^{\top}B\|\left(C(K^{(0)},\Sigma^{(0)})-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}}\leq\frac{1}{\|R\|+\gamma\frac{\|B^{\top}B\|\left(C\left(K^{(t)},\Sigma^{(t)}\right)-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}}$
[…]
and in the same way one could get $a\leq\frac{\tau}{2\left\|R+\gamma B^{\top}P_{K^{(t)}}B\right\|}$. Thus, Lemma 4.4 can be applied such that
$C\left(K^{(t+1)},\Sigma^{(t+1)}\right)-C(K^{*},\Sigma^{*})\leq(1-\zeta)\left(C\left(K^{(t)},\Sigma^{(t)}\right)-C(K^{*},\Sigma^{*})\right).$
and $aI\preceq\Sigma^{(t+1)}\preceq I.$ Thus, the induction is complete. Finally, observe that $0<\zeta\leq\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|}=\frac{\mu\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|r_{0}}<1,\text{ and }\zeta\leq\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}=\frac{\tau^{2}\sigma_{\min}(R)}{8r_{0}^{3}},$ the proof is complete.∎
@@ 5 Iterative Policy Optimization Method
By Bellman equation for the value function $J_{K,\Sigma}$,
@@
$K^{\prime},\Sigma^{\prime}=\arg\min_{\widetilde{K},\widetilde{\Sigma}}\mathbb{E}_{u\sim\pi_{\widetilde{K},\widetilde{\Sigma}}}\Big{[}x^{\top}Qx+u^{\top}Ru+\tau\log\pi_{\widetilde{K},\widetilde{\Sigma}}(u|x)+\gamma{J_{K,\Sigma}}(Ax+Bu+w)\Big{]}$
@@
$K^{(t+1)}$ $=K^{(t)}-\left(R+\gamma B^{\top}P_{K^{(t)}}B\right)^{-1}E_{K^{(t)}},$ (IPO)
@@ 5.1 Global Linear Convergence
$N\geq\frac{\left\|S_{K^{*},\Sigma^{*}}\right\|}{\mu}\log\frac{C\left(K^{(0)},\Sigma^{(0)}\right)-C\left(K^{*},\Sigma^{*}\right)}{\varepsilon},$
@@
Suppose $(K^{\prime},\Sigma^{\prime})$ follows a one-step updating rule of (IPO) from $(K,\Sigma)$:
$K^{\prime}$ $=K-\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K},\quad\Sigma^{\prime}=\frac{\tau}{2}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}.$ (25)
Then
@@
Theorem 5.1 suggests that (IPO) achieves a global linear convergence. Compared with (RPG), (IPO) exhibits faster convergence in terms of the rate at which the objective function decreases (cf. Lemmas 4.4 and 5.2). Furthermore, the subsequent section demonstrates that (IPO) enjoys a super-linear local convergence when the initial policy parameter is within a neighborhood of the optimal policy parameter.
5.2 Super-linear Local Convergence
This section establishes a super-linear local convergence for (IPO). We first introduce some constants used throughout this section:
$\xi_{\gamma,\rho}$ $:=\frac{1-\gamma\rho^{2}+\gamma}{(1-\gamma\rho^{2})^{2}},\qquad\zeta_{\gamma,\rho}:=\frac{2-\rho^{2}}{(1-\rho^{2})^{2}(1-\gamma)}+\frac{1}{(1-\rho^{2})^{2}(1-\gamma\rho^{2})},$ (26)
$\omega_{\gamma,\rho}$ $:=\frac{1}{(1-\rho^{2})(1-\gamma)}+\frac{1}{(1-\rho^{2})(1-\gamma\rho^{2})}.$

### v4

@@ 3 Analysis of Value Function and Policy Gradient
A function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is considered smooth if the following condition is satisfied: $\left|f(x)-f(y)+\nabla f(x)^{\top}(y-x)\right|\leq\frac{m}{2}\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{n},$ with $m$ a finite constant [15, 21]. In general, characterizing the smoothness of $C(K,\Sigma)$ is challenging, as it may become unbounded when the eigenvalues of ABKA-BK exceed $\frac{1}{\sqrt{\gamma}}$ or when $\sigma_{\min}(\Sigma)$ is close to 0. Nevertheless, in Lemma 3.3, we will see that if $C(K,\Sigma)$ is “almost” smooth, then the difference $C(K,\Sigma)-C(K^{\prime},\Sigma^{\prime})$ can be bounded by the sum of linear and quadratic terms involving $K-K^{\prime}$ and $\Sigma-\Sigma^{\prime}$.
@@ 4 Regularized Policy Gradient Method
RPG takes into account the inherent structure of the parameter space, which can accelerate convergence. By the explicit expressions of $\nabla_{K}C(K,\Sigma)$ and $\nabla_{\Sigma}C(K,\Sigma)$ in (3.6), the above update is equivalent to
(RPG) $K^{(t+1)}$ $=K^{(t)}-2\eta_{1}E_{K^{(t)}},$
@@
Assume that the covariance matrix $\Sigma$ is parameterized as scalar multiplication of an identity matrix, i.e., $\Sigma=sI$ for some $s>0$ and $\pi_{K,s}(x,u)=\mathcal{N}(Kx,sI).$ Then the natural policy gradient follows the update [29]:
(4.1) $K^{\prime}$ $=K-\eta G_{K}^{-1}\nabla C(K,sI),\,\,s^{\prime}=s-\eta G_{s}^{-1}\partial_{s}C(K,sI),$
where $G_{K}$ and $G_{s}$ are the Fischer information matrices under policy $\pi_{K,s}$, i.e.,
$G_{K}=\mathbb{E}\left[\sum_{t=0}^{\infty}\,\nabla\log\pi_{K,s}(u_{t}|x_{t})\nabla\log\pi_{K,s}(u_{t}|x_{t})^{\top}\right],$
$G_{s}=\mathbb{E}\left[\sum_{t=0}^{\infty}\,\partial_{s}\log\pi_{K,s}(u_{t}|x_{t})\partial_{s}\log\pi_{K,s}(u_{t}|x_{t})^{\top}\right].$
When the covariance matrix of the Gaussian policy takes a diagonal form as in $\pi_{K,s}$, (4.1) are equivalent to
(4.2) $K^{\prime}=K-\eta\nabla C(K,sI)(S_{K,sI})^{-1},\quad s^{\prime}=s-\tilde{\eta}~\partial_{s}C(K,sI)s^{2},$
@@
Even though (RPG) is similar to (4.2), there is no corresponding Fisher information form associated with (RPG) because of the additional step that simultaneously updates the covariance matrix $\Sigma$, which may not necessarily be diagonal.
@@
Given $\tau\in(0,2\sigma_{\min}(R)],$ take $(K^{(0)},\Sigma^{(0)})\in\Omega$ such that $\Sigma^{(0)}\preceq I$. Define $M_{\tau}:=\frac{\tau k}{2(1-\gamma)}\log\left(\frac{\sigma_{\min}(R)}{\pi\tau}\right)$ and $r_{0}:=\max\left\{\frac{2}{\tau\sigma_{\min}\left(\Sigma^{(0)}\right)},\,\|R\|+\gamma\frac{\|B^{\top}B\|\left(C(K^{(0)},\Sigma^{(0)})-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}\right\}.$ Then for $\eta_{1},\eta_{2}=\frac{1}{2r_{0}},\frac{\tau(1-\gamma)}{2r_{0}^{2}}$ and for $N\geq\max\left\{\frac{\|S_{K^{*},\Sigma^{*}}\|r_{0}}{\mu\sigma_{\min}(R)},\frac{8r_{0}^{3}}{\tau^{2}\sigma_{\min}(R)}\right\}\log\frac{C(K^{(0)},\Sigma^{(0)})-C(K^{*},\Sigma^{*})}{\varepsilon},$ the regularized policy gradient descent (RPG) has the following performance bound:
@@
Remark 4.3.
Theorem 4.2 shows that in order to achieve an $\epsilon$-optimal value function, the number of iterations required is at least $\mathcal{O}\left(\frac{1}{\tau^{5}}\right)$. Thus, the larger the value of regularization $\tau$, the smaller the number of iterations, and the faster the convergence.
Remark 4.4.
In the RPG update, we only need an upper bound and a lower bound for $\Sigma$, namely, $aI\preceq\Sigma\preceq bI$ for some $0<a<b.$ Different choices of $b$ may lead to different (admissible) ranges for $\tau$. For ease of exposition, we set $b=1$ in Theorem 4.2, and the results can be easily extended to the general case of any $b\geq a>0$.
To prove Theorem 4.2, we will first need the boundedness of the one-step update of the covariance $\Sigma$, in order to guarantee the well-definedness of the objective function along the trajectory when performing (RPG) (Lemma 4.5). Additionally, we will bound the one-step update of (RPG) (Lemma 4.6), and provide an upper bound of $\|P_{K}\|$ in terms of the objective function $C(K,\Sigma)$ (Lemma 4.7).
@@
Let $(K,\Sigma)\in\Omega$ be given such that $0\prec\Sigma\preceq I$. Fix $\tau\in(0,\,2\sigma_{\min}(R)]$, $a\in\Big(0,\,\min\Big\{\frac{\tau}{2\|R+\gamma B^{\top}P_{K}B\|},\sigma_{\min}(\Sigma)\Big\}\Big).$ Let $K^{\prime},\Sigma^{\prime}$ be the one-step update of $K,\Sigma$ using (RPG) with $\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}$. Then $aI\preceq\Sigma^{\prime}\preceq I$.
@@
with $0<\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}<1.$
@@
Let $M_{\tau}$ be defined in the same way as in Theorem 4.2. Then for all $(K,\Sigma)\in\Omega$, $C(K,\Sigma)\geq\left(\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)\right)\|P_{K}\|+M_{\tau}.$
Proof 4.8.
@@
(4.3) $\frac{1}{\left\|R+\gamma B^{\top}P_{K^{(t)}}B\right\|}\geq\frac{1}{\|R\|+\gamma\|B^{\top}B\|\left\|P_{K^{(t)}}\right\|}\geq\frac{1}{\|R\|+\gamma\frac{\|B^{\top}B\|\left(C({K^{(t)}},\Sigma^{(t)})-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}}.$
Let $a=\frac{\tau}{2r_{0}}$, $\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}$. The proof is completed by induction to show $C(K^{(t+1)},\Sigma^{(t+1)})\leq(1-\zeta)C(K^{(t)},\Sigma^{(t)})$ and $aI\preceq\Sigma^{(t+1)}\preceq I$ holds for all $t\geq 0$: at $t=0$, apply (4.3) to get $\eta_{1}\leq\frac{1}{2\|R+\gamma B^{\top}P_{K^{(0)}}B\|}$ and $a\leq\frac{\tau}{2\|R+\gamma B^{\top}P_{K^{(0)}}B\|}$. Additionally with $\eta_{2}=\frac{\tau(1-\gamma)}{2r_{0}^{2}}=\frac{2(1-\gamma)a^{2}}{\tau}$ and $aI\preceq\Sigma^{(0)}\preceq I$, Lemma 4.6 can be applied to get $C(K^{(1)},\Sigma^{(1)})\leq(1-\zeta)C(K^{(0)},\Sigma^{(0)})\leq C(K^{(0)},\Sigma^{(0)}),$ and $aI\preceq\Sigma^{(1)}\preceq I$. The proof proceeds by arguing that Lemma 4.6 can be applied at every step. If it were the case that $C\left(K^{(t)},\Sigma^{(t)}\right)\leq(1-\zeta)C\left(K^{(t-1)},\Sigma^{(t-1)}\right)\leq C\left(K^{(0)},\Sigma^{(0)}\right)$ and $aI\preceq\Sigma^{(t)}\preceq I$, then
$2\eta_{1}=\frac{1}{\|R\|+\gamma\frac{\|B^{\top}B\|\left(C(K^{(0)},\Sigma^{(0)})-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}}\leq\frac{1}{\|R\|+\gamma\frac{\|B^{\top}B\|\left(C\left(K^{(t)},\Sigma^{(t)}\right)-M_{\tau}\right)}{\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)}},$
thus by (4.3) $\eta_{1}\leq\frac{1}{\left\|R+\gamma B^{\top}P_{K^{(t)}}B\right\|}$ and in the same way $a\leq\frac{\tau}{2\left\|R+\gamma B^{\top}P_{K^{(t)}}B\right\|}$. Thus, Lemma 4.6 can be applied such that $C\left(K^{(t+1)},\Sigma^{(t+1)}\right)-C(K^{*},\Sigma^{*})\leq(1-\zeta)\left(C\left(K^{(t)},\Sigma^{(t)}\right)-C(K^{*},\Sigma^{*})\right),$ and $aI\preceq\Sigma^{(t+1)}\preceq I.$ Thus, the induction is complete. Finally, observe that $0<\zeta\leq\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|}=\frac{\mu\sigma_{\min}(R)}{\|S_{K^{*},\Sigma^{*}}\|r_{0}}<1,\text{ and }\zeta\leq\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}=\frac{\tau^{2}\sigma_{\min}(R)}{8r_{0}^{3}},$ the proof is complete.
@@ 5 Iterative Policy Optimization Method
By the Bellman equation for the value function $J_{K,\Sigma}$,
@@
$K^{\prime},\Sigma^{\prime}=\arg\min_{\widetilde{K},\widetilde{\Sigma}}\mathbb{E}_{u\sim\pi_{\widetilde{K},\widetilde{\Sigma}}}\Big[x^{\top}Qx+u^{\top}Ru+\tau\log\pi_{\widetilde{K},\widetilde{\Sigma}}(u|x)+\gamma{J_{K,\Sigma}}(Ax+Bu+w)\Big].$
@@
(IPO) $K^{(t+1)}$ $=K^{(t)}-\left(R+\gamma B^{\top}P_{K^{(t)}}B\right)^{-1}E_{K^{(t)}},$
@@ 5.1 Global Linear Convergence
$N\geq\frac{\left\|S_{K^{*},\Sigma^{*}}\right\|}{\mu}\log\frac{C(K^{(0)},\Sigma^{(0)})-C(K^{*},\Sigma^{*})}{\varepsilon},$
@@
Suppose $(K^{\prime},\Sigma^{\prime})$ follows a one-step updating rule of (IPO) from $(K,\Sigma)$. Then
@@
Theorem 5.1 suggests that (IPO) achieves a global linear convergence. Compared with (RPG), (IPO) exhibits faster convergence in terms of the rate at which the objective function decreases (cf. Lemmas 4.6 and 5.2). Furthermore, the subsequent section demonstrates that (IPO) enjoys a local super-linear convergence when the initial policy parameter is within a neighborhood of the optimal policy parameter.
5.2 Local Super-linear Convergence
This section establishes a local super-linear convergence for (IPO). We first introduce some constants used throughout this section:
(5.1) $\xi_{\gamma,\rho}$ $=\frac{1-\gamma\rho^{2}+\gamma}{(1-\gamma\rho^{2})^{2}},\qquad\zeta_{\gamma,\rho}=\frac{2-\rho^{2}}{(1-\rho^{2})^{2}(1-\gamma)}+\frac{1}{(1-\rho^{2})^{2}(1-\gamma\rho^{2})},$
$\omega_{\gamma,\rho}$ $=\frac{1}{(1-\rho^{2})(1-\gamma)}+\frac{1}{(1-\rho^{2})(1-\gamma\rho^{2})}.$